OpenStudy (anonymous):

4x + 10y = 94 y = 3x – 28 Answer a. (11, 5) b. (5, -11) c. (-11, 5) d. (5, 11) whats da anwers to dis

OpenStudy (anonymous):

the answer is a....just substitute x and y for both and you can see it checks out

OpenStudy (anonymous):

4(11)+10(5)=94 44+50=94 94=94 5=3(11)-28 5=33-28 5=5